# The Baldwin Effect in Advancing Generalizability of Physics-Informed Neural Networks

### v2

@@ III-D Baldwinian neuroevolution algorithm and hyperparameters
Our preliminary experiments show that the performance of Baldwinian neuroevolution is robust across a range of CMA-ES hyperparameters such as population size and initial standard deviation (std.) of search distribution, and the number of tasks sampled for fitness evaluation per iteration. Hence, a robust setting that shows good convergence in fitness across different types of problems is chosen based on initial experiments. In the majority of this study, we employ the CME-ES for evolving $(\theta,\lambda)$, although other results in Suppl. LABEL:subsec:additional_evo_depth show the extensibility of this framework to other neuroevolution algorithms.
In Algorithm 2, a fixed set of collocation points are used for the construction of least squares matrix and vector $(\mathbf{A},\mathbf{b})$ across the training tasks for ease of implementation. We choose the collocation points to coincide with the location of labelled data for $l_{\_}{MSE}$ computation, although this is not a prerequisite. Similarly, the same number of nonlinear iterations $N$ is used across nonlinear training tasks sampled from the task distribution for ease of implementation.
@@
In our preliminary experiments, we found it is critical to have the $l_{\_}{MSE}$ component in the optimization objective even though the learning of future test tasks remains solely physics-based. We set $\tau_{\_}{LSE}=\tau_{\_}{MSE}=1$ as default for the computation of overall fitness unless there is a huge difference in magnitude between $l_{\_}{LSE}$ and $l_{\_}{MSE}$ during Baldwinian neuroevolution. In addition, we can multiply the BC/IC rows in both $\mathbf{A}$ and $\mathbf{b}$ to re-balance the importance between PDE and BC/IC errors in the least-squares solution.
@@
Although this study focuses on the paradigm of Baldwinian neuroevolution as a pathway towards generalizable neural physics solvers, other combinations of neuroevolution or lifetime learning algorithms may be used. One key advantage of neuroevolution is that the fitness evaluations (population size $n_{\_}{pop}$$\times$ number of random tasks $n_{\_}{task}$) required each iteration can be easily parallelized across multiple GPUs to fully harvest any hardware advantage. In particular, we utilized the JAX framework to harness previously reported performance improvements for automatic differentiation and linear algebra operations [43, 44]. The experimental study is performed on a workstation with an Intel Xeon W-2275 Processor and 2 NVIDIA GeForce RTX 3090.
@@ IV Experimental Studies
We examine the efficacy of Baldwinian neuroevolution for learning physics (Algorithm 1 and Algorithm 2 in Section III-D) as formalized in a two-stage stochastic programming problem in  III-A. Several ODE/PDE problems which are representative of real-world phenomena are used to demonstrate Baldwinian neuroevolution for physics in the following sections. Table I summarizes various neuroevolution and Baldwinian-PINN lifetime learning configurations and performance on their respective test tasks.
@@ V Conclusion
Acknowledgment
This research is supported by A*STAR under the AME Programmatic project: Explainable Physics-based AI for Engineering Modelling & Design (ePAI) [Award No. A20H5b0142].

### v3

@@ III-D2 Evolutionary algorithm
In the majority of this study, we employ the covariance matrix adaptation evolution strategy (CMA-ES) [39] for evolving $(\theta,\lambda)$, although results in Suppl. LABEL:subsec:additional_evo_depth show the extensibility of this framework to other neuroevolution algorithms. As an instantiation of information-geometric optimization algorithms [40], CMA-ES represents the population of $(\theta,\lambda)$ in $\bm{\mathcal{P}}$ using a multivariate normal search distribution, initialized with zero mean and standard deviation (std.) as tuning hyperparameter. It iteratively adapts the search distribution based on the rank-based fitness landscape until the convergence criteria is met. Our experiments show that the performance of Baldwinian neuroevolution is robust across a range of CMA-ES hyperparameters such as population size and initial standard deviation (std.) of search distribution, and the number of tasks sampled for fitness evaluation per iteration. Hence, a robust setting that shows good convergence in fitness across different types of problems is chosen based on initial experiments.
The network weights and biases can take any value from (($\infty$( - ∞, $\infty$)∞ ), hence there is no restriction to the continuous search space of $\theta$ representing their distributional means and spreads. The learning hyperparameter $\lambda\geq 0$ can be evolved in continuous search space and its absolute value is then used for computing the least-squares solution. In our implementation with CMA-ES, $(\theta,\lambda)$ share the same initial standard deviation, and we scale the learning hyperparameter by a factor, i.e., $\lambda\leftarrow 1\mathrm{e}{-4}\times\text{abs}(\lambda)$, to improve the performance of the pseudoinverse.
@@
In our preliminary experiments, we found it helpful to have the $l_{\_}{MSE}$ component in the optimization objective even though the learning of future test tasks remains solely physics-based. We set $\tau_{\_}{LSE}=\tau_{\_}{MSE}=1$ as default for the computation of overall fitness unless there is a huge difference in magnitude between $l_{\_}{LSE}$ and $l_{\_}{MSE}$ during Baldwinian neuroevolution. In addition, we can multiply the BC/IC rows in both $\mathbf{A}$ and $\mathbf{b}$ to re-balance the importance between PDE and BC/IC errors in the least-squares solution.
@@
As this study focuses on the paradigm of Baldwinian neuroevolution as a pathway towards generalizable neural physics solvers, other combinations of evolutionary optimization or lifetime learning algorithms may be used. One key advantage of evolutionary optimization is that the fitness evaluations (population size $n_{\_}{pop}$$\times$ number of random tasks $n_{\_}{task}$) required each iteration can be easily parallelized across multiple GPUs to fully harvest any hardware advantage. In particular, we utilized the JAX framework to harness previously reported performance improvements for automatic differentiation and linear algebra operations [41, 42]. The experimental study is performed on a workstation with an Intel Xeon W-2275 Processor and 2 NVIDIA GeForce RTX 3090.
@@ IV Experimental Studies
We examine the efficacy of Baldwinian neuroevolution for learning physics (Algorithm 1 and Algorithm 2 in Section III-B) as formalized in a two-stage stochastic programming problem in Section III-A. Several ODE/PDE problems which are representative of real-world phenomena are used to demonstrate Baldwinian neuroevolution for physics in the following sections. Table I summarizes various neuroevolution and Baldwinian-PINN lifetime learning configurations and performance on their respective test tasks.